OpenStudy (destinyyyy):

Help with this question please ~ What mass of H2So4 must be dissolved to 2.40 kg H2O to produce a 1.20m solution?

OpenStudy (destinyyyy):

What i have so far is 1.20m = mol H2So4/ 2.40 kg H20 I have no clue what your meaning. Its already in mol form so why would i need to calculate h2so4?

thomaster (thomaster):

ow lol is misread sorry XD give me a sec

OpenStudy (destinyyyy):

lol ok just confused me for a sec there

thomaster (thomaster):

you know you need to make a solution of 1.2 mol per liter H2O You have 2.4 KG H2O, when no density of temperature is given you can assume that 2.4 kg water equals 2.4 liter. Per liter H2O you need 1.2 mol H2SO4 How many mol H2SO4 do you need for 2.4 liter?

OpenStudy (destinyyyy):

What??? O.o

thomaster (thomaster):

2.4 kg H20 = 2.4 L H20 for 1 liter H20 you need 1.2 mol H2SO4 how many mol H2SO4 do you need for 2.4 liter To solve this just multiply 2.4 liter by 1.2 mol. because: 1 liter is 1 * 1.2 mol 2.4 liter is 2.4 * 1.2 mol = 2.88 mol So you need to dissoolve 2.88 mol H2SO4 in 2.4 liter H2O to get a 1.2M solution (M= molarity = mol per liter) Last thing you need to do is convert mol H2SO4 to gram H2SO4 To do this you need to find the molecular weight of H2SO4. To find this you need a periodic table and look what's the weight of H (hydrogen), S (sulfur) and O (oxygen). write down these 3 weights (they are displayed in gram/mol) then count the amount of H, S and O you have. in this case that's 2x H. 1x S and 4x O. so 2x the mass of 1 hydrogen atom and 1x sulfur etc then sum them up and you have the molecular weight.

thomaster (thomaster):

when you know the molecular weight of H2SO4, you know what 1 mole of H2SO4 weights. You need to know how much 2,88 mol weights so just multiply the molecular weight of H2SO4 by 2.88

OpenStudy (destinyyyy):

Why change it to liters? Its already in mol form

OpenStudy (destinyyyy):

I honestly do not understand this. I got how to solve for molality. But i dont understand how to do this. My teacher solved for kg solvent but for this problem i have to solve for mol solute. And i dont have an example problem to help figure it out.

thomaster (thomaster):

why change it to liters? Molarity does not work with weight, only volume.

thomaster (thomaster):

\[Molarity (M) = \frac{mol }{ Liter }\]

thomaster (thomaster):

H= 1.00794 g/mol S= 32.065 g/mol O= 15.9994 g/mol you have 2x H, 1x S and 4x O 2 * 1.00794 = 2.01588 1 * 32.065 = 32.065 4 * 15.9994 = 63.9976 2.01588+32.065+63.9976 = 98.07848 grams per mol 1 mol = 98.07848 gram 2.88 mol = 2.88 * 98.07848 = 282.46602 grams That's your final answer You would need 282.46602 gram H2SO4 to dissolve in 2.40 kg H2O to produce a 1.20M solution.